In a class of 32 students, 18 offers chemistry, 16 offers Physics 22 offer Mathematics. 6 offer all three subjects, 3 offer Chemistry and Physics only and 5 offer only physics only. Each student offers at least one subject. Find the number of student who offer A. chemistry only B. only one student C. only two subjects


1
Expert's answer
2021-03-24T08:01:20-0400
"N(P\cup C\cup M)=32"

"N(P)=16, P(C)=18, P(M)=22"


"N(P\cap C\cap M)=6"

"N(P\cap C \cap M^C)=3"

"N(P\cap C^C \cap M^C)=5"

"N(P\cap M \cap C^C)=N(P)-N(P\cap C\cap M)-"

"-N(P\cap C \cap M^C)-N(P\cap C^C \cap M^C)"

"=16-6-3-5=2"

"N(P\cup C\cup M)=N(P)+N(M)+N(C)"

"-N(P\cap M \cap C^C)-N(P\cap C \cap M^C)-"

"-N(C\cap M \cap P^C)-2N(P\cap C\cap M)"

"N(C\cap M \cap P^C)=16+18+22"

"-32-3-2-2(6)=7"


A. "N(C\cap P^C\cap M^C)=2"


B. "N(\text{only one subject})=5+2+7=14"


C. "N(\text{only two subjects})=3+2+7=12"